# Neural Ordinary Differential Equations for Mapping the Magnetic QCD Phase Diagram via Holography

### v3

@@ Abstract
The QCD phase diagram is crucial for understanding strongly interacting matter under extreme conditions, with major implications for cosmology, neutron stars, and heavy-ion collisions. We present a novel holographic QCD model utilizing neural ordinary differential equations (ODEs) to map the QCD phase diagram under magnetic field $B$, baryon chemical potential $\mu_{B}$, and temperature $T$. By solving the inverse problem of constructing the gravitational theory from Lattice QCD data, we reveal an unprecedentedly rich phase structure at finite $B$, including discovering multiple critical endpoints (CEPs) under strong magnetic fields. Specifically, for $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, we identify two distinct CEPs at $(T_{C}=87.3\,\mathrm{MeV},\,\mu_{C}=115.9\,\mathrm{MeV})$ and $(T_{C}=78.9\,\mathrm{MeV},\,\mu_{C}=244.0\,\mathrm{MeV})$. Notably, the critical exponents vary depending on the CEP’s location. These findings significantly advance our understanding of the QCD phase structure and provide concrete predictions for experimental validation at upcoming facilities such as FAIR, JPARC-HI, and NICA.
@@
Introduction—Quantum Chromodynamics (QCD), the theory of strong interactions, governs quark and gluon behavior under extreme conditions, underpinning high-energy physics, astrophysics, and cosmology. The QCD phase diagram, which maps transitions between the deconfined quark-gluon plasma and hadronic matter, is central to understanding phenomena in non-central heavy-ion collisions, neutron stars, and the early universe Braun-Munzinger:2008szb ; Philipsen:2012nu ; Gupta:2011wh . The global structure of the QCD phase diagram, spanning temperature ($T$), baryon chemical potential ($\mu_{B}$), and magnetic field ($B$), remains poorly understood Andersen:2014xxa . At low $\mu_{B}$ and $B=0$, QCD matter undergoes a smooth crossover, while at high $\mu_{B}$, a first-order phase transition occurs. Locating the critical endpoint (CEP) separating these regimes is a primary goal of heavy-ion collision experiments Aoki:2006we ; HotQCD:2018pds .
Magnetic fields, generated in non-central heavy-ion collisions and present in astrophysical objects like magnetars, significantly influence the QCD equation of state. Strong magnetic fields modify the critical temperature, induce novel phases, and play a key role in shaping the QCD phase structure Andersen:2014xxa , making $B$ an essential factor in searching CEP experimentally STAR:2020tga STAR:2021fge STAR:2022hbp . Exploring the magnetic QCD phase structure remains a significant challenge.
@@
With substantial progress, current non-perturbative approaches to studying the QCD phase diagram still face significant challenges. As a first-principle approach, Lattice QCD encounters the well-known sign problem at finite baryon density, making numerical simulations at high $\mu_{B}$ computationally prohibitive Philipsen:2012nu . Moreover, simulating high magnetic field configurations demands substantial computational resources, leading to limited exploration of the QCD phase structure in these regions Andersen:2014xxa . These limitations significantly hinder our understanding of the interplay among $T$, $\mu_{B}$, and $B$, particularly in the regions of the phase diagram where experimental observations are most challenging.
@@
Neural ordinary differential equations (neural ODEs) excel in modeling complex, continuous dynamic systems by optimizing over infinite-dimensional parameter spaces with adaptive precision. We leverage neural ODEs in holography to address these challenges, offering a data-driven mechanism that surpasses traditional machine learning methods relying on empirical formulas for finite-dimensional parameter fitting Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2020jug ; Hashimoto:2022eij ; Chen:2024ckb . This innovative approach captures the intricate dependencies of thermodynamic quantities on $T$, $\mu_{B}$, and $B$ with unprecedented accuracy and flexibility footnote .
These advantages allow our approach to a concrete global structure of the phase diagram and uncover unprecedented novel features of the QCD phase structure, such as multiple CEPs and their associated critical behaviors, which were previously never identified. In particular, for sufficiently large $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, the $T$-$\mu_{B}$ plane exhibits two CEPs, a feature exceeding prior expectations. Furthermore, we analyze the critical behavior near the CEPs, finding that the critical exponents satisfy scaling relations but vary based on the CEP’s location. This innovative combination of machine learning and holography significantly advances our understanding of the QCD phase structure and provides new avenues for better agreement with experimental and lattice QCD results.
Holographic model–To capture essential QCD dynamics at finite magnetic field, temperature, and baryon chemical potential, we employ a holographic framework based on the five-dimensional gravitational theory:
@@
where $\kappa_{N}^{2}$ is the effective Newton constant. The metric $g_{\mu\nu}$ characterizes spacetime geometry, and the real scalar field $\phi$ accounts for conformal symmetry breaking. The Maxwell field $A_{\mu}$ with $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ introduces a finite baryon number density, while the magnetic field $B$ is described by another Maxwell field $\hat{A}_{\mu}$ with $\hat{F}_{\mu\nu}=\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$. The functions $Z(\phi)$, $\hat{Z}(\phi)$, and $V(\phi)$ that encode the non-perturbative features of our system are calibrated against lattice QCD data.
@@
The magnetic field $B$ breaks Lorentz invariance along the $z$-axis, leading to anisotropic pressure. Solving the bulk equations of motion (EoMs) allows us to extract thermodynamic quantities, such as the free energy density $\Omega$, longitudinal pressure $P_{z}$, entropy density $s$, baryon density $n_{B}$, and magnetization $M$. For further details, see Supplementary Material Supplementary-Material . This holographic approach provides a first-principles description of QCD dynamics under extreme conditions.
Figure 1: Schematic of the neural ODE approach for solving the inverse problem in QCD phase diagram analysis. Initial conditions from the black hole (BH) horizon are connected to QCD boundary data via EoMs. The neural network initialized as a trial function for the coupling $\hat{Z}(\phi)$ is optimized via backpropagation to minimize the loss function L, representing the difference from lattice QCD data. Parameters $\xi$ are optimized using gradient descent. Technical details are provided in the Supplementary Material Supplementary-Material .
@@
Figure 2: Thermodynamic Quantities from Holographic QCD Model vs. Lattice Data. Temperature dependence of (a) magnetic susceptibility $\chi_{B}$, (b) magnetization $M$, (c) entropy density $s/T^{3}$, and (d) longitudinal pressure $\Delta p_{z}=p_{z}|_{B}-p_{z}|_{B=0}$ across magnetic fields. Shaded regions show lattice QCD estimates Bali:2014kia ; solid lines indicate model predictions. Here, $e=1$, giving $B=1\,\text{GeV}^{2}=1.602\times 10^{19}$ Gauss.
@@
Neural ODEs—In the absence of a first-principle method to determine the coupling functions in our bottom-up model, we constrain these functionals using available lattice QCD data—a challenge known as the inverse problem. As lattice QCD data at finite $B$ and $\mu_{B}$ become more abundant, manually tuning control parameters becomes impractical. Instead of traditional neural networks that approximate solutions in a fixed finite-dimensional parameter space Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2020jug ; Hashimoto:2022eij ; Chen:2024ckb , our neural ODE approach systematically explores the infinite-dimensional parameter space of the coupling functions, achieving higher precision and accuracy.
@@
In practice, we impose boundary conditions at the ultraviolet (UV) boundary and the black hole event horizon. By solving the EoMs numerically, the neural ODE generates a trial equation of state that is iteratively optimized through backpropagation to reduce deviations from lattice QCD data. This process efficiently converges to an optimal magnetic coupling, providing a good agreement with lattice data and refining model predictions across unexplored magnetic fields and chemical potentials (see Fig. 1 for illustration). A detailed algorithmic breakdown of this methodology is included in the Supplementary Material Supplementary-Material . We construct a (2+1)-flavor holographic QCD (hQCD) model using this neural ODE approach to achieve a precise fit with lattice QCD results Bali:2014kia . A similar neural ODE-based method is applied to determine $Z(\phi)$ and $V(\phi)$ by matching to data at $B=0$ HotQCD:2014kol ; Borsanyi:2021sxv . The resulting functional forms are benchmarked rigorously and agree with recent lattice simulations and experimental data Cai:2022omk ; Li:2023mpv .
@@
Fig. 2 presents our holographic predictions for four independent thermodynamic quantities: magnetic susceptibility $\chi^{B}$, magnetization $M$, entropy density $s$, and longitudinal pressure $p_{z}$, compared to lattice QCD data Bali:2014kia . We find good agreement across the available magnetic fields, supporting our holographic model. This is the first holographic model to achieve good agreement with lattice data for magnetic fields up to $B=0.6\text{ GeV}^{2}$. Further analysis is provided in the Supplementary Material, showing consistency with lattice QCD results Bali:2014kia .
@@
QCD phase diagram—With the model fully established, we construct the QCD phase diagram at finite $B$, $T$, and $\mu_{B}$ by computing the free energy density $\Omega$. The full phase diagram is depicted in Fig. 3. The light blue area denotes the first-order phase transition surface, dividing the quadrant into two parts: the high-temperature region corresponds to the quark-gluon plasma, while the low-temperature region corresponds to the hadron gas phase. The deep blue line in the diagram marks the location of CEP for various magnetic fields, where the first-order phase transition terminates and transitions into a smooth crossover at small chemical potentials. At $B=0$, the phase diagram was presented in Fig. 3 of Cai:2022omk , where the first-order transition line terminates at $(T_{C}=105\,\text{MeV},\mu_{C}=555\,\text{MeV})$. Similarly, the phase structure in the $T$-$B$ plane at $\mu_{B}=0$ reveals a line of first-order transitions ending at the CEP with $(T_{C}=89.6\,\text{MeV},B=1.6\,\text{GeV}^{2})$, consistent with lattice QCD predictions Cuteri:2023evl .
Fig. 3 highlights the following three key observations longversion .
@@
    As the magnetic field $B$ increases up to $B=1.764\,\text{GeV}^{2}$, the critical chemical potential $\mu_{C}$ at the CEP decreases, indicating that stronger magnetic fields shift the CEP to lower chemical potentials. This shows a significant effect of magnetic fields on the CEP location, affecting conditions in experimental settings like heavy-ion collisions.
@@
    The critical temperature $T_{C}$ at the CEP initially decreases with increasing $B$, reaching a minimum before increasing again. This turning point occurs around $T=80\text{MeV}$, $B=1.6\sim 1.7\text{GeV}^{2}$, and $\mu_{B}=0.2\sim 0.28\text{GeV}$. This behavior suggests complicated effects in the presence of a background magnetic field. It could related to the inverse magnetic catalysis and magnetic catalysis reported in the literature.
@@
Critical exponents—Beyond mapping the phase diagram, we examine critical behavior near the CEPs via critical exponents. These exponents describe how thermodynamic quantities, such as susceptibility and specific heat, diverge near critical points, typically following power-law scaling. These exponents help identify the CEP’s universality class and provide insights into QCD transitions under extreme conditions.
@@
Table 1 presents the critical exponents for the CEP at three different magnetic field values, denoted by hQCD (I, II, III) longversion . The critical exponents satisfy the scaling relations $\alpha+2\beta+\gamma=2$ and $\alpha+\beta(1+\delta)=2$, ensuring the self-consistency of our results. Although close to mean-field values, these exponents show significant deviations depending on the CEP location, particularly as $B$ increases. Such deviations highlight the features of our holographic QCD model, which cannot be attributed to large-$N$ effects typical in conventional holographic duality, where mean-field behavior is expected. Similar deviation happened in the holographic 2-flavor model Zhao:2023gur , the critical exponents match those of the quantum 3D Ising model, further emphasizing the distinct nature of our approach compared to traditional large-$N$ QCD models.
@@
Table 1: Critical exponents from experiments in non-QCD fluids, the full quantum 3D Ising model, mean-field (van der Waals) theory, the DGR model DeWolfe:2010he , and our (2+1)-flavor hQCD model. The hQCD (I, II, III) correspond to the critical exponents for $\mu_{B}=554.66\ \text{MeV},B=0$ (hQCD I), $\mu_{B}=501.4\ \text{MeV},B=0.3\ \text{GeV}^{2}$ (hQCD II), and $\mu_{B}=0,B=1.6\ \text{GeV}^{2}$ (hQCD III), respectively.

### v4

@@ Abstract
The QCD phase diagram is crucial for understanding strongly interacting matter under extreme conditions, with major implications for cosmology, neutron stars, and heavy-ion collisions. We present a novel holographic QCD model utilizing neural ordinary differential equations (ODEs) to map the QCD phase diagram under magnetic field $B$, baryon chemical potential $\mu_{B}$, and temperature $T$. By solving the inverse problem of constructing a gravitational theory from Lattice QCD data, we reveal an unprecedentedly rich phase structure at finite $B$, including multiple critical endpoints (CEPs) in strong magnetic fields. Specifically, for $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, we identify two distinct CEPs at $(T_{C}=87.3\,\mathrm{MeV},\,\mu_{C}=115.9\,\mathrm{MeV})$ and $(T_{C}=78.9\,\mathrm{MeV},\,\mu_{C}=244.0\,\mathrm{MeV})$. Notably, the critical exponents vary depending on the CEP’s location, and the conventional scaling relations can be violated in the presence of strong magnetic fields. These findings significantly advance our understanding of the QCD phase structure and provide concrete predictions for experimental validation at upcoming facilities such as FAIR, JPARC-HI, and NICA.
Keywords: Gauge-gravity duality, QCD phase transition, Critical exponents, Machine learning
I I. Introduction
@@
Quantum Chromodynamics (QCD), the theory of strong interactions, governs quark and gluon behavior under extreme conditions, underpinning high-energy physics, astrophysics, and cosmology. The QCD phase diagram, which maps transitions between the deconfined quark-gluon plasma and hadronic matter, is central to understanding phenomena in non-central heavy-ion collisions, neutron stars, and the early universe Braun-Munzinger:2008szb ; Philipsen:2012nu ; Gupta:2011wh . The global structure of the QCD phase diagram, spanning temperature ($T$), baryon chemical potential ($\mu_{B}$), and magnetic field ($B$), remains poorly understood Andersen:2014xxa . At low $\mu_{B}$ and $B=0$, QCD matter undergoes a smooth crossover, while at high $\mu_{B}$, a first-order phase transition occurs. Locating the critical endpoint (CEP) separating these regimes is a primary goal of heavy-ion collision experiments Aoki:2006we ; HotQCD:2018pds . Magnetic fields, generated in non-central heavy-ion collisions and present in astrophysical objects like magnetars, significantly influence the QCD equation of state. Strong magnetic fields modify the critical temperature, induce novel phases, and play a key role in shaping the QCD phase structure Andersen:2014xxa , making $B$ an essential factor in searching for CEP experimentally STAR:2020tga ; STAR:2021fge ; STAR:2022hbp . Exploring the magnetic QCD phase structure remains a significant challenge.
@@
With substantial progress, current non-perturbative approaches to studying the QCD phase diagram still face significant challenges. As a first-principle approach, Lattice QCD encounters the well-known sign problem at finite baryon density, making numerical simulations at high $\mu_{B}$ computationally prohibitive Philipsen:2012nu . Moreover, simulating high-magnetic-field configurations requires substantial computational resources, limiting exploration of the QCD phase structure in these regions Andersen:2014xxa . These limitations significantly hinder our understanding of the interplay among $T$, $\mu_{B}$, and $B$, particularly in the regions of the phase diagram where experimental observations are most challenging.
@@
Neural ordinary differential equations (neural ODEs) excel in modeling complex, continuous dynamic systems by optimizing over infinite-dimensional parameter spaces with adaptive precision. In this work, we leverage neural ODEs in holography to address these challenges, offering a data-driven mechanism that surpasses traditional machine learning methods Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2020jug ; Hashimoto:2022eij ; Chen:2024ckb . This innovative approach captures the intricate dependencies of thermodynamic quantities on $T$, $\mu_{B}$, and $B$ with unprecedented accuracy and flexibility.
These advantages allow our approach to yield a concrete global structure of the phase diagram and uncover unprecedented features of the QCD phase structure, such as multiple CEPs and their associated critical behaviors, which were previously unobserved. In particular, for sufficiently large $B=1.618\,\mathrm{GeV}^{2}=2.592\times 10^{19}$ Gauss, the $T$-$\mu_{B}$ plane exhibits two CEPs, a feature exceeding prior expectations. Furthermore, we analyze the critical behavior near the CEPs and find that the critical exponents vary depending on the location of the CEP. Notably, the conventional scaling relations can be violated in the presence of strong magnetic fields. This innovative combination of machine learning and holography significantly advances our understanding of the QCD phase structure and provides new avenues for better agreement with experimental and lattice QCD results.
This work is organized as follows. We introduce the holographic model in Section II. We then describe our method for fixing the bulk theory using machine learning in Section III. The full QCD phase diagram and its key features are summarized in Sections IV and V. We conclude and provide future directions in Section VI. In Appendix A, we additionally present details for solving the equations of motion and the definitions of thermodynamic quantities. A detailed algorithmic breakdown of the neural ODEs is included in Appendix B.
Figure 1: Schematic of the neural ODE approach for solving the inverse problem in QCD phase diagram analysis. Initial conditions from the black hole (BH) horizon are connected to QCD boundary data via EoMs. The neural network initialized as a trial function for the coupling $\hat{Z}(\phi)$ is optimized via backpropagation to minimize the loss function L, representing the difference from lattice QCD data. Parameters $\xi$ are optimized using gradient descent. Technical details are provided in Appendix B.
II II. Holographic model
To capture essential QCD dynamics at finite magnetic field, temperature, and baryon chemical potential, we employ a holographic framework based on the five-dimensional gravitational theory:
@@
where $\kappa_{N}^{2}$ is the effective Newton constant. The metric $g_{\mu\nu}$ characterizes spacetime geometry, and the real scalar field $\phi$ accounts for conformal symmetry breaking. The Maxwell field $A_{\mu}$ with $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ introduces a finite baryon number density, while the magnetic field $B$ is described by another Maxwell field $\hat{A}_{\mu}$ with $\hat{F}_{\mu\nu}=\partial_{\mu}\hat{A}_{\nu}-\partial_{\nu}\hat{A}_{\mu}$. The functions $Z(\phi)$, $\hat{Z}(\phi)$, and $V(\phi)$ that encode the non-perturbative features of our system are calibrated against lattice QCD data. The form of $V(\phi)$ and $Z(\phi)$ is taken from Cai:2022omk .
[…]
$Z(\phi)$ $={1\over 1+c_{3}}{\text{sech}[c_{4}\phi^{3}]}+{c_{3}\over 1+c_{3}}e^{-c_{5}\phi}\,.$
Note that the effective number of model parameters cannot be reduced by coordinate or gauge transformations. Coordinate transformations simplify the metric but leave $V(\phi)$ and $Z(\phi)$ unchanged, where the parameters $c_{1},c_{2},c_{3},c_{4},c_{5}$ govern the scalar field dynamics (see Appendix A). Similarly, gauge transformations simplify the gauge fields without affecting the parameters in $V(\phi)$ or $Z(\phi)$, which remain independent of the gauge choice. The comparison of the holographic model (2) to experimental observations can be found in Li:2023mpv ; He:2023ado .
By varying the action (II), the equations of motion (EoMs) can be obtained:
$\nabla_{\mu}\nabla^{\mu}\phi-{\partial_{\phi}Z\over 4}F_{\mu\nu}F^{\mu\nu}-{\partial_{\phi}\hat{Z}\over 4}\hat{F}_{\mu\nu}\hat{F}^{\mu\nu}-\partial_{\phi}V=0\,$, (3)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\nabla^{\nu}(ZF_{\nu\mu})=0\,$,
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\nabla^{\nu}(\hat{Z}\hat{F}_{\nu\mu})=0\,$,
$R_{\mu\nu}-{1\over 2}Rg_{\mu\nu}={1\over 2}\nabla_{\mu}\phi\nabla_{\nu}\phi+{Z\over 2}F_{\mu\rho}{F_{\nu}}^{\rho}+{\hat{Z}\over 2}\hat{F}_{\mu\rho}{\hat{F}_{\nu}}^{\rho}$
$+{1\over 2}\left(-{1\over 2}\nabla_{\mu}\phi\nabla^{\mu}\phi-{Z\over 4}F_{\mu\nu}F^{\mu\nu}-{\hat{Z}\over 4}\hat{F}_{\mu\nu}\hat{F}^{\mu\nu}-V\right)g_{\mu\nu}\,.$
The bulk black hole solutions in Anti-de Sitter spacetime (AdS) are
$ds^{2}=-f(r)e^{-\eta(r)}dt^{2}+{dr^{2}\over f(r)}+r^{2}(dx^{2}+dy^{2}+g(r)dz^{2})\,$, (4)
$\phi=\phi(r),\;A_{\mu}dx^{\mu}=A(r)dt,\;\hat{A}_{\mu}dx^{\mu}={B\over 2}(xdy-ydx)\,$,
where $r$ is the holographic radial coordinate with the AdS boundary located at $r\rightarrow\infty$. The source term $\phi_{s}\equiv\lim_{r\rightarrow\infty}r\phi(r)$ essentially breaks the conformal symmetry and plays the role of the energy scale. Substituting the above ansatz into (3) yields the bulk equations that will be solved numerically.
The blackening function $f(r)$ is vanishing at the event horizon $r=r_{h}$. The temperature and entropy density of the QCD system are given by the famous Hawking temperature and Bekenstein-Hawking entropy of a black hole:
$T={1\over 4\pi}f^{\prime}(r_{h})e^{-\eta(r_{h})/2},\quad s={2\pi\over\kappa_{N}^{2}}r_{h}^{3}\,$, (5)
@@
evaluated at the black hole event horizon. Note that the magnetic field $B$ breaks Lorentz invariance along the $z$-axis, leading to anisotropic pressure. Solving the bulk EoMs allows us to extract thermodynamic quantities via the holographic renormalization, such as the free energy density $\Omega$, energy density $\epsilon$, longitudinal pressure $p_{z}$, baryon density $n_{B}$, and magnetization $M$. Moreover, they are found to satisfy the expected thermodynamic laws:
$\begin{split}Ts&=\epsilon-\Omega-\mu_{B}n_{B}\,,\\ d\Omega&=-sdT-n_{B}d\mu_{B}-MdB\,.\end{split}$ (6)
For further details, see Appendix A. This holographic approach provides a phenomenological description of QCD dynamics under extreme conditions. We have used units $c=\hbar=k_{B}=e=1$.
Figure 2: The magnetic coupling $\hat{Z}(\phi)$ as a function of $\phi$ from machine learning. The black solid curve is the one obtained from our neural ODEs architecture, and the red dotted one from the fitting function (S23) in Appendix B.
@@
Figure 3: Thermodynamic Quantities from Holographic QCD Model vs. Lattice Data. Temperature dependence of (a) magnetic susceptibility $\chi_{B}$, (b) magnetization $M$, (c) entropy density $s/T^{3}$, and (d) longitudinal pressure $\Delta p_{z}=p_{z}|_{B}-p_{z}|_{B=0}$ across magnetic fields. The shaded regions show lattice QCD estimates Bali:2014kia ; the solid lines indicate model predictions. Here, $e=1$, giving $B=1\,\text{GeV}^{2}=1.602\times 10^{19}$ Gauss.
III III. Neural ODEs
@@
In the absence of a first-principle method to determine the coupling functions in our bottom-up model, we constrain these functionals using available lattice QCD data—a challenge known as the inverse problem. As lattice QCD data at finite $B$ and $\mu_{B}$ become more abundant, manually tuning control parameters becomes impractical. Our neural ODE approach systematically explores the infinite-dimensional parameter space of the coupling functions, achieving higher precision and accuracy.
@@
In practice, we impose boundary conditions at the ultraviolet (UV) boundary and the black hole event horizon. By solving the EoMs numerically, the neural ODE generates a trial equation of state, which is then iteratively optimized via backpropagation to reduce deviations from lattice QCD data. This process efficiently converges to an optimal magnetic coupling, providing a good agreement with lattice data and refining model predictions across unexplored magnetic fields and chemical potentials (see Fig. 1 for illustration). A detailed algorithmic breakdown of this methodology is included in Appendix B. We construct a (2+1)-flavor holographic QCD (hQCD) model using this neural ODE approach to achieve a precise fit with lattice QCD results Bali:2014kia . A similar neural ODE-based method is applied to determine $Z(\phi)$ and $V(\phi)$ by matching to data at $B=0$ HotQCD:2014kol ; Borsanyi:2021sxv . The resulting functional forms are benchmarked rigorously and agree with recent lattice simulations and experimental data Cai:2022omk ; Li:2023mpv . The profile for $\hat{Z}(\phi)$ from our neural ODEs is presented in Fig. 2. This strong non-monotonicity highlights the necessity for parameter tuning in machine learning.
@@
Fig. 3 presents our holographic predictions for four independent thermodynamic quantities: magnetic susceptibility $\chi^{B}$, magnetization $M$, entropy density $s$, and longitudinal pressure $p_{z}$, compared to lattice QCD data Bali:2014kia . We find good agreement across the available magnetic fields, supporting our holographic model. This is the first holographic model to achieve good agreement with lattice data for magnetic fields up to $B=0.6\text{ GeV}^{2}$. Further analysis is provided in Appendix B, showing consistency with lattice QCD results Bali:2014kia . While Fig. 3 demonstrates an overall reasonable agreement with the lattice data, especially for the entropy density, there are deviations from the central value, which can be as large as $2\sigma$, particularly at finite magnetic fields. These discrepancies are due to several factors, including the inherent limitations of the neural ODE approach and the simplifications made in the model. The neural ODE framework, while powerful for data-driven optimization, has certain limitations in capturing the full complexity of the system. This includes challenges related to the parameter space and the trade-off between model complexity and fitting accuracy. Further refinement of the model and optimization process could help reduce these deviations in future work.
Figure 4: The phase diagram on the $B$-$T$ plane at vanishing $\mu_{B}$. The blue dot denotes the CEP, and the blue line corresponds to the first-order line.
IV IV. QCD phase diagram
@@
With the model fully established, we construct the QCD phase diagram at finite $B$, $T$, and $\mu_{B}$ by computing the free energy density $\Omega$. The full phase diagram is depicted in Fig. 5. The light blue area denotes the first-order phase transition surface, dividing the quadrant into two parts: the high-temperature region corresponds to the quark-gluon plasma, while the low-temperature region corresponds to the hadron gas phase. The deep blue line in the diagram marks the location of CEP for various magnetic fields, where the first-order phase transition terminates and transitions into a smooth crossover at small chemical potentials. At $B=0$, the phase diagram was presented in Fig. 3 of Cai:2022omk , where the first-order transition line terminates at $(T_{C}=105\,\text{MeV},\mu_{C}=555\,\text{MeV})$. Fig. 4 shows the phase structure in the $T$-$B$ plane at $\mu_{B}=0$, revealing a line of first-order transitions ending at the CEP located at $(T_{C}=89.6\,\text{MeV},B=1.6\,\text{GeV}^{2})$, consistent with lattice QCD predictions Cuteri:2023evl .
Fig. 5 highlights the following three key observations.
@@
    As the magnetic field $B$ increases up to $B=1.764\,\text{GeV}^{2}$, the critical chemical potential $\mu_{C}$ at the CEP decreases, indicating that stronger magnetic fields shift the CEP to lower chemical potentials. This shows a significant effect of magnetic fields on the CEP location, which affects conditions in experimental settings such as heavy-ion collisions.
@@
    The critical temperature $T_{C}$ at the CEP initially decreases with increasing $B$, reaching a minimum before increasing again. This turning point occurs around $T=80\text{MeV}$, $B=1.6\sim 1.7\text{GeV}^{2}$, and $\mu_{B}=0.2\sim 0.28\text{GeV}$. This behavior suggests complicated effects in the presence of a background magnetic field. It could be related to the inverse magnetic catalysis and magnetic catalysis reported in the literature.
@@
V V. Critical exponents
Beyond mapping the phase diagram, we examine critical behavior near the CEPs via critical exponents. These exponents describe how thermodynamic quantities, such as susceptibility and specific heat, diverge near critical points, typically following power-law scaling. These exponents help identify the CEP’s universality class and provide insights into QCD transitions under extreme conditions.
@@
Figure 7: Critical phenomena of the CEP in the $B$-$T$ plane at $\mu_{B}=0$. Top left: Specific heat $C_{n}$ along the tangent of the first-order line, yielding the critical exponent $\alpha=0.00917$. Top right: Discontinuity of the entropy density $s$ across the first-order phase transition line, from which the critical exponent $\beta=0.3944$ is extracted. Bottom left: Baryon susceptibility along the first-order axis, which gives the critical exponent $\gamma=1.05715$. Bottom right: Entropy density $s$ as a function of $\mu_{B}$ on the critical isotherm $T=T_{\text{CEP}}$, yielding the critical exponent $\delta=3.9878$.
For illustration, we now present the critical behavior of the CEP in the $B$-$T$ plane at zero $\mu_{B}$ (see Fig. 4). For example, the discontinuity in entropy density across the first-order phase transition line is displayed in the top-right panel of Fig. 7, while the bottom-right panel shows the trajectory of entropy density $s$ as $\mu_{B}$ approaches $\mu_{B\text{CEP}}$ on the critical isotherm. The power-law scaling is evident in both cases, yielding the critical exponents $\beta=0.3944$ and $\delta=3.9878$. The values of these two critical exponents deviate clearly from the mean-field predictions. Similarly, we obtain the other critical exponents and $\alpha=0.00917$ and $\gamma=1.05175$ from the left panels of Fig. 7. Table 1 presents the critical exponents for the CEP at three different magnetic field values, denoted by CEP (I, II, III). Although close to mean-field values, these exponents show significant deviations depending on the CEP location, particularly as $B$ increases. Such deviations highlight the features of our holographic QCD model, which cannot be attributed to large-$N$ effects typical in conventional holographic duality, where mean-field behavior is expected. A similar deviation was observed in the holographic 2-flavor model Zhao:2023gur . The critical exponents match those of the quantum 3D Ising model, further emphasizing the distinct nature of our approach compared to traditional large-$N$ QCD models. The gradual change in critical exponents observed at these CEPs is unusual in standard critical phenomena, where transitions between universality classes are typically abrupt. However, this gradual variation in our model arises from the magnetic field’s influence on the system’s symmetry and degrees of freedom. In particular, the magnetic field reduces certain degrees of freedom, modifying the system’s overall symmetry, which in turn causes the observed gradual transition in critical behavior.
@@
Table 1: Critical exponents from experiments in non-QCD fluids, the full quantum 3D Ising model, mean-field (van der Waals) theory, the DGR model DeWolfe:2010he , and our (2+1)-flavor hQCD model. The CEP (I, II, III) correspond to the critical exponents for $\mu_{B}=554.66\ \text{MeV},B=0$ (CEP I), $\mu_{B}=501.4\ \text{MeV},B=0.3\ \text{GeV}^{2}$ (CEP II), and $\mu_{B}=0,B=1.6\ \text{GeV}^{2}$ (CEP III), respectively.
Before ending this section, we point out that the four critical exponents $(\alpha,\beta,\gamma,\delta)$ are not independent by assuming that the singular part of the free energy density near a CEP is a generalized homogeneous function, from which one obtains two scaling relations:
$\alpha+2\beta+\gamma=2,\quad\delta-\gamma/\beta=1\,.$ (7)
The first is the Rushbrooke relation, and the second is the Widom relation. At a weak magnetic field, within numerical uncertainty, the critical exponents we obtain are quantitatively consistent with both scaling relations, as seen for CEP I and CEP II in Table 1. However, a clear violation of the scaling relations emerges at a strong magnetic field. For CEP III in Table 1 (see also Fig. 7), we find that
$\alpha+2\beta+\gamma=1.8498,\quad\delta-\gamma/\beta=1.3212\,$,
deviating from the conventional scaling laws in (7). These departures can be attributed to several factors. First, numerical uncertainties in the model’s calculations can cause small deviations in the exponents. Second, the presence of a strong magnetic field may affect the scaling behavior, leading to departures from the traditional relations. Finally, the violation may reflect intrinsic properties of the holographic model, where quantum fluctuations and the competition among multiple degrees of freedom in strongly correlated systems could break the standard scaling hypothesis. Although violations of (7) are uncommon, similar deviations have been reported in other black hole systems, see e.g. Frassino:2014pha ; Hu:2024ldp .
VI VI. Conclusion